The larger number of two numbers is 7 less than 3 times the smaller number . The sum of two numbers is 37. Find the two numbers
Answer by ikleyn(52784)

.
x + (3x-7) = 37


4x - 7 = 37


4x = 37 + 7 = 44


x = 11  is the smaller number.


Find the larger number on your own.

Good luck and happy learning !